# People who frequently use ChatGPT for writing tasks are accurate and robust detectors of AI-generated text

### v1

@@ 3 Fine-grained analysis of expert performance
Table 4: Taxonomy of clues used by experts to explain their detection decisions. For each category, we report the frequency of explanations that mention that category (regardless of if the annotator was correct) and provide examples of explanations for both human-written and AI-generated articles. While vocabulary and sentence structure form the most frequent clues, more complex phenomena like originality, clarity, formality, and factuality are also distinguishing features.
The previous section establishes that human experts are highly accurate detectors of AI-generated text even in the presence of evasion attempts. Here, we first provide in-depth comparisons between experts and automatic detectors. Next, we perform a fine-grained coding of expert explanations into different categories (e.g., vocabulary, originality, formality), which allows us to examine what they focus on during the different experiments. Finally, we analyze differences between annotators and what they focus on when they are incorrect, highlighting implications for the future training of human annotators for AI-generated text detection.
@@
We begin with a direct comparison between our human experts and state-of-the-art AI detectors. Specifically, we run five different detectors on our set of 300 articles:
The middle rows of Table 3 show that only Pangram Humanizers (average TPR of 99.3% with FPR of 2.7% for base model) matches the human expert majority vote, and it also outperforms each expert individually. Pangram is near perfect on the first four experiments and falters just slightly on humanized o1-Pro articles, while GPTZero struggles significantly on o1-Pro with and without humanization. The open-source detectors degrade in the presence of paraphrasing and underperform both closed detectors by large margins on average. We again note that our experts are untrained at this detection task, and they could likely improve their individual performance if provided with feedback on their errors.
@@
Our automatic baselines are largely unexplainable: they either rely on combinations of various statistical properties of the text, or operate as opaque classifiers trained on large datasets. In contrast, we can easily solicit explanations from human annotators, and in this section we use GPT-4o to code these explanations into a schema (Table 4) developed by the authors after careful manual analysis. Details of the coding process used to label explanations can be found in §D.
Vocabulary-related clues are mentioned in the majority (53.1%) of all explanations, while sentence structure (35.9%), grammar (24.8%), and originality (23.7%) are also common. We note that many of these categories (e.g., originality, factuality, tone) are much more difficult to assess automatically than others (e.g., vocabulary), and these may currently be areas where humans have an advantage over automatic detectors.
Figure 3: (Top) A heatmap displaying the frequency with which annotators mentioned specific categories in their explanations when they were correct. Interestingly, vocabulary becomes a less frequent clue for o1-Pro-generated articles, especially with humanization. (Bottom) Same as above, except only computed over explanations when experts were incorrect. Formality is a big source of misdirection for o1-Pro articles, while fixating on sentence structure can lead experts to false positives. Details of each category can be found in Table 4.
@@
Figure 3 (upper) shows the frequency that clue categories are mentioned in explanations for which the expert makes the correct decision. We observe several interesting shifts as the complexity of the article generation setup increases: for example, 57.1% of correct explanations about o1-Pro articles mention vocabulary, compared to only 42.3% for the humanized o1-Pro articles prompted to avoid “AI vocab”. Somewhat counterintuitively, explanations about paraphrased articles note AI vocab in 88% of explanations, compared to only 69.8% of non-paraphrased GPT-4o articles. Similarly, quotations are mentioned in 33.8% of explanations about paraphrased articles, a much higher rate than other configurations: a close reading of explanations reveals that experts flagged quotes that were always in the same format and style (e.g., only placed at the end of each paragraph). Originality is a common clue for detecting human-written text, highlighting the gulf in creativity between humans and LLMs Chakrabarty et al. (2024).
@@
Our experts focus on different properties of the text to arrive at their decisions. Annotator 1 is the only one to pick up on “AI names”: in fact, 63.3% of GPT-4o and 70% of Claude-3.5-Sonnet articles include either the name Emily or Sarah.121212o1-Pro favors names of real people instead of fictional names when generating articles. Annotator 2 and 3 emphasize linguistic features, such as humans not always following “proper” writing conventions or the tendency of LLMs to always list examples in groupings of three. Annotator 4 focuses more on the flow of the articles, analyzing the specificity of detail and motivation behind an article, while Annotator 5 frequently mentions whether a quotation sounds natural or not. Further analysis of individual performance can be found in §D.2. This diversity partly explains why an ensemble of expert annotators performs so well, and it also suggests room for training so that all experts are at least aware of most distinguishing features.
@@
Figure 3 (lower) shows the frequency that clue categories are mentioned in explanations for which experts make incorrect decisions. Despite near-perfect majority vote performance, experts make several errors individually, particularly with the qualitative change in text produced by o1-Pro. For Annotator 3, who relies heavily on linguistic features, o1-Pro’s frequent use of contractions and colloquialisms dramatically lowered their detection rate: they mention formality in none of their explanations for Experiments 2 and 3, compared to 66.7% of o1-Pro and 83.3% of humanized o1-Pro explanations as shown in Figure 11. Studying expert false positives is also insightful Figure 3 (lower): 31% of explanations for false positives mention vocabulary, typically when human-written content contains words like “delve” and “crucial” that our experts associate with AI-generated text. More prominently, 50% of false positives focus on sentence structure, when humans write stylistically-similar sentences to those preferred by LLMs (e.g., contiguous blocks of sentences with similar lengths, multiple lists of three items).
@@ 4 Can LLMs be prompted to mimic human expert detectors?
Our experiments so far highlight the benefits of hiring humans to perform AI-generated text detection: in aggregate, they are more accurate and robust than all automatic detectors we tried. Additionally, they can provide detailed explanations of their decision-making process, unlike all of the automatic detectors in our study. However, an obvious drawback is that hiring humans is expensive and slow: on average, we paid \$2.82 per article including bonuses, and we gave annotators roughly a week to complete a batch of 60 articles. In this section, we prompt LLMs to imitate our expert annotators, providing the guidebook used in our humanization experiments (§2.5) to the model and asking it to produce an explanation and a label. While the approach shows promise, outperforming detectors such as Binoculars and RADAR, it fails to reach the performance of our human experts as well as advanced automatic detectors like Pangram.
@@
Using our guidebook for AI detection (Table 11), we prompt an off-the-shelf LLM to decide whether a candidate text is human-written or AI-generated and explain its decision based on the criteria in the guidebook (see Table 21 for prompt template). While we can control this detector’s behavior by modifying the prompt (e.g., provide explanations before producing a final label), it does not output a single score, which makes setting a false positive threshold impossible. [^13] We do not experiment with thresholding the token-level probabilities of the labels “AI” vs. “Human”, since these are inaccessible with closed LLMs, but this is a promising route for future research on open prompt-based detectors. We derive the final output of our detector by using the prediction from the detection model in a deterministic configuration.
Prompt-based detection shows promise but struggles with humanization and high false positive rates:
We implement our prompt-based detector by zero-shot prompting GPT-4o-2024-11-20 and o1, [^14] We set temperature=0 for GPT-4o and reasoning efforts to ‘medium’ for o1 (note that OpenAI has fixed temperature=1 for o1). Further experiments can be found in Table 22. comparing the effects of chain-of-thought (CoT) prompting Wei et al. (2022) and the inclusion of the detection guide on its performance. For GPT-4o, the best detector configuration is achieved with both CoT and the guidebook (average TPR 78%), while o1 is best with CoT but without the guidebook (54%). We observe that o1 is more conservative than GPT-4o, demonstrating lower FPRs on average at the cost of lower TPR. However, it is also more capable at detecting humanized text. On average, our best configuration with GPT-4o-2024-11-20 performs comparably to Binoculars and Fast-DetectGPT, although obviously at a much higher cost. The gap between this method and humans demonstrates that we have a long way to go in order to “teach” LLMs how to detect AI-generated text like humans do; we speculate that fine-tuning may be helpful to close the gap.
@@ 5 Related work
Our work builds on prior research centering around AI-generated text detection using both human and automatic approaches, as well as recent work on evading such detectors.
@@
Our study most closely resembles several papers published prior to ChatGPT’s release. Existing work notes that while naïve annotators do not reliably detect AI-generated texts Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021); Karpinska et al. (2021), some annotators manage to perform very well Ippolito et al. (2020). More recently, these findings have been reproduced in other domains such as Italian news articles Puccetti et al. (2024) and poetry Porter and Machery (2024). Dugan et al. (2020) and Dugan et al. (2022) address a more complex task where humans and machines are tasked to identify boundary between human and AI-generated text in mixed documents.
Automatic detection of AI-generated text:
Successful automatic detection methods are typically either perplexity-based Mitchell et al. (2023); Bao et al. (2023); Hans et al. (2024) or trained classifiers Solaiman et al. (2019); Emi and Spero (2024); Verma et al. (2023). Some detectors have focused on sentence-level detection Kushnareva et al. (2024); Wang et al. (2023), emphasizing the need for more detailed and explainable detection. Dugan et al. (2024), Li et al. (2024), and Zhang et al. (2024) attempt to compare the performance of detectors in different domains and adversarial attack settings. However, most automatic detectors are unreliable in the face of attacks such as paraphrasing Krishna et al. (2023); Sadasivan et al. (2024) and highly sensitive to stylistic factors Doughman et al. (2025).
Evading detection:
Increasing LLM usage among the general population has spurred corresponding interest in humanization of AI-generated text. [^15] Relevant discussions on Reddit: ”AI Humanizer Recommendations?” (2024), ”How to humanize ai-generated texts?” (2024), ”How to humanize the AI generated content?” (2024) In fact, many commercial humanization services exist to fill this demand. [^16] Some humanization services include Undetectable AI, MyEssayWriter.ai and Stealth Writer. Initial research has proposed several humanization attacks to evade automatic detection Wang et al. (2024b); Lu et al. (2023); Shi et al. (2023); Wang et al. (2024a), while Zhou et al. (2024) and Masrour et al. (2025) improve the robustness of automatic detectors to humanized content.
@@
Many frameworks group errors in machine-generated text into categories similar to the ones we use in this paper Gehrmann et al. (2019); Dou et al. (2021). More recently, Ma et al. (2023) discover gaps in terms of depth and content quality between scientific AI-generated and human-written text. Shaib et al. (2024) introduces syntactic templates, noting the repetitiveness of AI-generated text in comparison to human-written references. More closely related to our work is that of Ji et al. (2024), who qualitatively categorize human detection explanations but do not employ expert annotators. Jakesch et al. (2023) explores the heuristics humans use to detect AI-generated text, with some similar conclusions to ours (e.g., humans are likely to associate more informal text with human writing).
@@ 6 Conclusion
Our paper demonstrates that a population of “expert” annotators—those who frequently use LLMs for writing-related tasks—are highly accurate and robust detectors of AI-generated text without any additional training. The majority vote of five such experts performs near perfectly on a dataset of 300 articles, outperforming all automatic detectors except the commercial Pangram model (which the experts match). Analysis of explanations provided by our expert annotators reveals that they pick up on not just vocabulary and sentence structure-related clues but also more complex properties like originality, factuality, and tone. We observe that each expert organically focuses on different aspects of the text, and we conjecture that with explicit training, human annotators can be made even more robust to advances in LLMs as well as evasion tactics (e.g., humanization). Future work can also explore human annotators working alongside automatic detectors like Pangram to improve detection accuracy and explainability. We find it apt to end with a comment from one of our experts about a particularly formulaic conclusion generated by GPT-4o:
@@
Our study is limited to articles in American English, chosen for their consistent formatting and high quality (i.e., professionally written and proofread). We also did not investigate factual accuracy, as it did not appear to be a significant cue for our annotators, who covered a broad range of topics. Finally, while we selected articles from reputable sources, there remains a possibility that some included AI-generated edits beyond our scope of detection.
@@ Appendix A Human Evaluation
Figure 6: Interface for annotators, with an example annotation from Annotator #4 with a humanized article from §2.5. This is the same article displayed in Figure 1. An annotator can highlight texts, make their decision, put confidence, and write an explanation. This AI-generated article was based off of In Alaska, a pilot drops turkeys to rural homes for Thanksgiving, written by Mark Thiessen & Becky Bohrer, was originally published by Associated Press on November 28, 2024.
@@ A.1 Finding Expert Annotators
The annotations for understanding the limitations of humans to detect AI-generated texts were done by 5 annotators who passed performance requirements based on the article expiriment of subsection 2.1. Only one annotator from the original experiment met these requirements. We recruited 10 more native English speakers on Upwork to take a 5 question sample of the original article task, questions which at most 2 out of 5 of the original annotators got correct. These 10 recruited annotators all had experience editing LLM generated content, frequently used LLMs, and had a professional background in editing or writing. Those who got at least 4 out of 5 correct (80%) given the rest of the 60 articles, and annotators who got at least 90% correct (54/60) were considered to be experts and recruited for the rest of the labeling rounds. 5 out of 10 annotators passed the 5 question trial; out of those 5, 4 passed the 60 question set.
@@ B.2 AI Text Generation
We generate the articles by prompting the models with the articles title, subtitle, approximate length, publication, and section. For stories, we prompt models with the title of the reddit thread. All closed-source models were prompted using the provider’s API,  [^23] The estimated cost for generating all data with each model is as follows: GPT-4o \$1.85USD, o1-Pro \$3.81 USD, Claude-3.5-Sonnet \$0.51 USD. All models were prompted and articles with the prompt presented in Table 8. Instructions for rounding to an approximate length were included to ensure that articles on the same topic would be of similar length. Statistics on the distribution of lengths by trial are presented in Figure 8
@@
We use the sentence-level paraphrasing approach outlined in PostMark Chang et al. (2024), changing the exact language in the prompts slightly for our use case. This approach paraphrases the text on a sentence by sentence level. For the initial sentence, only that sentence is given to be paraphrased. For all following sentences, the portion of the text already paraphrased is passed to the LLM, intending to improve the overall flow of the paraphrased article. We use GPT-4o as the paraphraser model with a temperature set to 0. The prompt for the initial sentence can be found in Table 9 and the prompt for all following sentences can be found in Table 10.
@@
Table 11: A truncated version of the AI Text Detection Guide. The full guide is located at https://github.com/jenna-russell/human_detectors.
@@
Table 14: The one article the majority of annotators did not detect correctly, generated from o1-Pro as part of Experiment 4 (see §2.4). "New Telescope Could Potentially Identify Planet X" was originally written by Emilie Le Beau Lucchesi in Discover Magazine on Nov. 6th, 2024

### v2

@@ 3 Analysis of expert performance
Table 3: Truncated taxonomy of clues used by experts to explain their detection decisions (see Table 17 for full version). For each category, we report the frequency of explanations that mention that category (regardless of correctness) and provide sample explanations for human-written and AI-generated articles. While vocabulary is the most common clue, complex phenomena like originality, quotes and clarity are also distinguishing features.
@@
We compare human experts with five AI detectors, including two closed-source classifiers, Pangram Emi and Spero (2024); Masrour et al. (2025) and GPTZero Tian and Cui (2023), and three open-source methods: Binoculars Hans et al. (2024), Fast-DetectGPT Bao et al. (2023), and RADAR Hu et al. (2023). [^20] We report TPR and FPR of automatic detectors using recommended thresholds whenever possible. For Fast-DetectGPT, we use a threshold calibrated to an FPR of 5% on a held-out set of 40 human-written articles. Thresholds used are found in §E.1. We evaluate Pangram in both its base and humanized modes, ignoring the “Possibly AI” label. Table 2 shows that only Pangram Humanizers (average TPR of 99.3% with FPR of 2.7% for base model) matches the human expert majority vote, and it also outperforms each expert individually, faltering just slightly on humanized o1-Pro articles. GPTZero struggles significantly on o1-Pro with and without humanization. Open-source detectors degrade in the presence of paraphrasing and underperform both closed detectors. We again note that our experts are untrained at this detection task, and they could likely improve their individual performance if provided with feedback.
Figure 3: A heatmap displaying the frequency with which annotators mentioned specific categories in their explanations when they were correct. Interestingly, vocabulary becomes a less frequent clue for o1-Pro-generated articles, especially with humanization. A heat map of corresponding incorrect explanation is displayed in Figure 9. Details of each category can be found in Table 3.
@@
In contrast to opaque automatic detectors, we can easily solicit explanations from human annotators. We use GPT-4o to code these explanations into a schema (Table 3) developed by the authors after careful manual analysis. Details of the coding process used to label explanations can be found in §D. Vocabulary-related clues are mentioned in the majority (53.1%) of all explanations, while sentence structure (35.9%), grammar (24.8%), and originality (23.7%) are also common. We note that many of these categories (e.g., originality, factuality, tone) are much more difficult to assess automatically than others (e.g., vocabulary), and these may currently be areas where humans have an advantage over automatic detectors.
@@
Figure 3 shows the frequency that clue categories are mentioned in explanations for which the expert makes the correct decision. We observe several shifts as the complexity of the article generation setup increases: for example, 57.1% of correct explanations about o1-Pro articles mention vocabulary, compared to only 42.3% for the humanized o1-Pro articles prompted to avoid “AI vocab”. Somewhat counterintuitively, explanations about paraphrased articles note AI vocab in 88% of explanations, compared to only 69.8% of non-paraphrased GPT-4o articles. Quotations are mentioned in a surprisingly high 33.8% of explanations about paraphrased articles: a close reading reveals that experts flagged quotes that were always in the same format and style (e.g., only placed at the end of each paragraph). Finally, the frequency of originality clues highlights the gulf in creativity between humans and LLMs Chakrabarty et al. (2024).
@@
Our experts focus on different properties of the text to arrive at their decisions, as described further in §D.3. Annotator 1 is the only one to pick up on “AI names”, while in fact, 63.3% of GPT-4o and 70% of Claude-3.5-Sonnet articles include either the name Emily or Sarah. 212121o1-Pro favors names of real people instead of fictional names when generating articles. Annotators 2 and 3 emphasize grammaticality or the tendency of LLMs to always list examples in groupings of three. Annotator 4 focuses on article flow, analyzing the specificity of detail and motivation behind an article, and Annotator 5 examines how natural any quotations sound. This diversity explains why an ensemble of expert annotators performs so well and it also suggests room for training so that all experts are at least aware of most distinguishing features. implies additional training can improve individual experts.
@@
Figure 9 shows clue frequencies in incorrect expert explanations (see §D.3 for more). For example, o1-Pro’s frequent use of contractions and colloquialisms fooled Annotator 3: they mention formality in none of their explanations for Experiments 2 and 3, compared to 66.7% of o1-Pro and 83.3% of humanized o1-Pro explanations (Figure 12). Studying expert false positives is also insightful (Figure 3): 31% of explanations here mention vocabulary, typically when human-written content contains “AI vocab” like delve and crucial.
When is it worth it to use human detection?
Human experts deliver detailed, freeform explanations and confidence ratings - features that automatic detectors generally lack, as they usually provide only a scalar score or binary label, limiting their trustworthiness in critical settings Lipton (2018). For example, before accusing a student of plagiarism or flagging a high-stakes legal document, someone may want to point to specific properties of the text that raised suspicion, or ensure that a false positive won’t have severe consequences. In our study, the majority vote of five expert annotators—without any specialized training—misclassified only 1 out of 300 articles, on par with the most accurate commercial detector (Pangram), even under adversarial paraphrasing and humanization. However, manual review cannot scale to high-volume, low-stakes tasks such as online forum moderation, where automated systems remain the only practical option.
@@ 4 Can LLMs be prompted to mimic human expert detectors?
We have established that human experts are more accurate, robust, and interpretable than automatic detectors. Can we prompt LLMs to mimic the decision-making process of human experts? In this section, we develop a prompt-based detector that also generates explanations, using the guidebook from §2.5. While the approach promisingly outperforms Binoculars and RADAR, it lags human experts and closed-source detectors like Pangram.
@@
We prompt both GPT-4o and o1 to decide whether a candidate text is human-written or AI-generated and explain its decision based on the criteria in the guidebook (see Table 22 for prompt template). [^22] We set temperature=0 for GPT-4o and reasoning efforts to ‘medium’ for o1 (note that OpenAI has fixed temperature=1 for o1). Further experiments can be found in Table 23. While we can control this detector’s behavior via its prompt (e.g., with and without chain-of-thought), it produces the label text instead of a scalar score, making it impossible to set an FPR threshold.
Prompt-based detection shows promise but struggles with humanization and high FPRs:
For GPT-4o, the best detector configuration is achieved with both CoT and the guidebook (average TPR 78%), while o1 is best with CoT but without the guidebook (54%). We observe that o1 is more conservative than GPT-4o, demonstrating lower FPRs on average at the cost of lower TPR. Our best configuration with GPT-4o-2024-11-20 performs comparably to Binoculars and Fast-DetectGPT, although at a much higher cost. We hypothesize that the gap between prompt-based detection and humans can be reduced by fine-tuning.
@@ 5 Related work
Research conducted prior to ChatGPT’s release concludes that while naïve annotators do not reliably detect AI-generated texts Ippolito et al. (2020); Brown et al. (2020); Clark et al. (2021); Karpinska et al. (2021), some annotators perform very well Ippolito et al. (2020); Dugan et al. (2022). Other research explores detection in Porter and Machery (2024) and other languages Puccetti et al. (2024); Wang et al. (2025). Dugan et al. (2020) and Dugan et al. (2022) address a more complex task of identifying the boundary between human and AI-generated text.
Automatic detection:
Successful automatic detection methods rely on either perplexity Mitchell et al. (2023); Bao et al. (2023); Hans et al. (2024) or trained classifiers Solaiman et al. (2019); Emi and Spero (2024); Verma et al. (2023). Prior work compares detectors’ performance in different domains and regimes Dugan et al. (2024); Li et al. (2024); Zhang et al. (2024). Detectors can be vulnerable to paraphrasing attacks Krishna et al. (2023); Sadasivan et al. (2024) and style changes Doughman et al. (2025). Studies have also examined humanization attacks to bypass detection Wang et al. (2024b); Lu et al. (2023); Shi et al. (2023); Wang et al. (2024a), while Zhou et al. (2024) and Masrour et al. (2025) boost detector robustness.
@@
Many frameworks group errors in AI-generated text similarly to our paper Gehrmann et al. (2019); Dou et al. (2021). Ma et al. (2023) discover gaps in depth and content quality between AI-generated and human-written scientific text, while Shaib et al. (2024) notes the repetitiveness of AI-generated text by introducing syntactic templates. Ji et al. (2024) categorize human nonexpert detection explanations.
@@ 6 Conclusion
Our paper demonstrates that a population of “expert” annotators—those who frequently use LLMs for writing-related tasks—are highly accurate and robust detectors of AI-generated text without any additional training. The majority vote of five such experts performs near perfectly on a dataset of 300 articles, outperforming all automatic detectors except the commercial Pangram model (which the experts match). Analysis of explanations provided by our expert annotators reveals that they pick up on not just vocabulary and sentence structure-related clues but also more complex properties like originality and tone. We observe that experts focus on different aspects of the text, and we conjecture that with explicit training, human annotators can be made even more robust to advances in LLMs as well as evasion tactics (e.g., humanization). Future work can also explore human annotators working alongside automatic detectors like Pangram to improve detection accuracy and explainability. We find it apt to end with a comment from one of our experts about a particularly formulaic conclusion generated by GPT-4o:
@@
Our study is limited to articles in American English, chosen for their consistent formatting and high quality (i.e., professionally written and proofread). We also did not investigate factual accuracy, as it did not appear to be a significant cue for our annotators, who covered a broad range of topics. While we selected articles from reputable sources, there remains a possibility that some included AI-generated edits beyond our scope of detection.
@@ Appendix A Human Evaluation
Study Design
: We employ a within-subjects design for our experiments, where each annotator judges both the human-written and AI-generated articles. This reduces variability from individual differences and requires fewer annotators Allen (2017). To further minimize bias, annotators are unaware of the pairing, and the article order is randomized.
@@
Figure 6: Interface for annotators, with an example annotation from Annotator #4 with a humanized article from §2.5. This is the same article displayed in Figure 1. An annotator can highlight texts, make their decision, put confidence, and write an explanation. This AI-generated article was based off of In Alaska, a pilot drops turkeys to rural homes for Thanksgiving, written by Mark Thiessen & Becky Bohrer, and was originally published by Associated Press on Nov. 28, 2024.
@@ A.1 Finding Expert Annotators
In our pilot study (§C), only the annotator who uses LLMs daily to edit AI-generated content achieved high accuracy. Prior work has noted on average, annotators cannot reliably detect AI-generated content Brown et al. (2020); Clark et al. (2021), but vary widely in ability Dugan et al. (2022); in this paper, we aim to investigate the upper bound of human performance. To find this limit, and understand the lexical clues humans use to recognize AI-generated content, we recruit four additional annotators with similar backgrounds to the high-performing annotator.
The annotations for understanding the limitations of humans to detect AI-generated texts were done by 5 annotators who passed performance requirements based on the article experiment of subsection 2.1. Only one annotator from the original experiment met these requirements. We recruited 10 more native English speakers on Upwork to take a 5 question sample of the original article task, questions which at most 2 out of 5 of the original annotators got correct. These 10 recruited annotators all had experience editing LLM generated content, frequently used LLMs, and had a professional background in editing or writing. Those who got at least 4 out of 5 correct (80%) given the rest of the 60 articles, and annotators who got at least 90% correct (54/60) were considered to be experts and recruited for the rest of the labeling rounds. 5 out of 10 annotators passed the 5 question trial; out of those 5, 4 passed the 60 question set.
While some individuals unfamiliar with LLMs may prove to be reliable detectors, we focus on the reliability of people who frequently use LLMs for writing tasks (e.g. English teachers, editors, publishers), as our pilot study suggests their expertise enables them to consistently identify systematic patterns characteristic of AI-generated text.
@@ B.1 Article Corpus
Exp # Model Human-written (n=30) AI-generated (n=30)
1 GPT-4o 652.5170.9 625.1126.9
2 Claude-3.5-Sonnet 793.1170.3 710.8133.5
3 GPT-4o paraphrased 777.8175.9 654.9110.6
4 o1-Pro 728.4155.1 813.8233.6
5 o1-Pro humanized 739.0128.3 815.6171.2
@@ B.2 AI Text Generation
We generate the articles by prompting the models with the articles title, subtitle, approximate length, publication, and section. For stories, we prompt models with the title of the reddit thread. All closed-source models were prompted using the provider’s API,  [^29] The estimated cost for generating all data with each model is as follows: GPT-4o \$1.85USD, o1-Pro \$3.81 USD, Claude-3.5-Sonnet \$0.51 USD. All models were prompted and articles with the prompt presented in Table 8. Instructions for rounding to an approximate length were included to ensure that articles on the same topic would be of similar length. Statistics on the distribution of lengths by trial are presented in Figure 8. An example article generation prompt is below:
Title: The Science That Makes Baseball Mud ‘Magical’
Article:
@@
We use the sentence-level paraphrasing approach outlined in PostMark Chang et al. (2024), changing the exact language in the prompts slightly for our use case. This approach paraphrases the text on a sentence by sentence level. For the initial sentence, only that sentence is given to be paraphrased. For all following sentences, the portion of the text already paraphrased is passed to the LLM, intending to improve the overall flow of the paraphrased article. Additionally, since our experts easily identified AI-generated named entities (e.g., Sarah Thompson) in the previous experiments, we extract named entities from the human-written article and instruct GPT-4o to include these names in its generated article. We use GPT-4o as the paraphraser model with a temperature set to 0. The prompt for the initial sentence can be found in Table 9 and the prompt for all following sentences can be found in Table 10.
@@
Table 11: A truncated version of the AI Text Detection Guide.
@@
Table 14: The one article the majority of annotators did not detect correctly, generated from o1-Pro as part of Experiment 4 (see §2.4). "New Telescope Could Potentially Identify Planet X" was originally written by Emilie Le Beau Lucchesi in Discover Magazine on Nov. 6th, 2024.
Additional Existing Humanizers
Recently, the general population is much more interested in evading automatic detectors,  [^30] Relevant discussions on Reddit: ”AI Humanizer Recommendations?” (2024), ”How to humanize ai-generated texts?” (2024), ”How to humanize the AI generated content?” (2024) many commercial humanization services exist to fill this demand.Some humanization services include Undetectable AI, MyEssayWriter.ai and Stealth Writer.